# Exact computation of Transfer Entropy with Path Weight Sampling

### v1

@@ Appendix F Variants of transfer entropy
$=\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{i}(k+1),X_{j}(k+1)|X_{j,[0,k]}\right)}{P\left(X_{i}(k+1)|X_{j,[0,k]}\right)P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (35)
which quantifies how much the prediction of $X_{i}(k+1)$ is improved by using $X_{j}(k+1)$ additional to the past trajectory $X_{j,[0,k]}$. Computation of $\widehat{\mathcal{T}}_{X_{i}\to X_{j}}$ requires marginalization over only one conditional distribution, $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$. Each of the probabilities in Eq. 35 can be computed by averaging analytically available transition probabilities over this conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, which is provided by TE-PWS. Thus, TE-PWS can be used to compute filtered transfer entropy at half the computational cost as Schreiber’s transfer entropy.

### v2

@@ Appendix G Variants of transfer entropy
$=\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{i}(k+1),X_{j}(k+1)|X_{j,[0,k]}\right)}{P\left(X_{i}(k+1)|X_{j,[0,k]}\right)P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (S33)
which quantifies how much the prediction of $X_{i}(k+1)$ is improved by using $X_{j}(k+1)$ in addition to the past trajectory $X_{j,[0,k]}$. The computation of $\widehat{\mathcal{T}}_{X_{i}\to X_{j}}$ requires marginalization over only one conditional distribution, $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$. Each of the probabilities in Eq. S33 can be computed by averaging analytically available transition probabilities $P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$, $P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ and $P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$, over this conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, which is provided by TE-PWS through the RR scheme. Thus, TE-PWS can be used to compute filtered transfer entropy at half the computational cost as Schreiber’s transfer entropy.